# Towards Sustainable Universal Deepfake Detection with Frequency-Domain Masking

### v1

@@ 5.1. Suggestions for Future Work
Several promising directions emerge from this work. Combining frequency masking with spatial augmentations or geometric transformations could yield complementary benefits, potentially capturing both local and global artifacts; in particular, translation appears synergistic with frequency masking, whereas rotation should be used judiciously. Developing self-supervised versions of this approach would reduce reliance on labeled fake data. The framework could be extended to video by incorporating temporal frequency analysis. There’s also significant potential in creating adaptive masking strategies that automatically adjust to different generator types and select transforms (e.g., FFT vs. DCT) based on data characteristics. These extensions would build upon the current findings while addressing the identified limitations, moving toward more comprehensive deepfake detection solutions.
@@ Acknowledgements.
This research is supported by the National Research Foundation, Singapore, under its AI Singapore Programmes (AISG Award No.: AISG2-TC-2022-007) and the SUTD project PIE-SGP-AI-2018-01, as well as by the Agency for Science, Technology and Research (A*STAR) under its MTC Programmatic Funds (Grant No. M23L7b0021). Additional support was provided by Changi General Hospital and the Singapore University of Technology and Design under the HealthTech Innovation Fund (HTIF Award No. CGH-SUTD-2021-004). The computational work was carried out partly using the resources from Sigma2—the National Infrastructure for High-Performance Computing and Data Storage in Norway, with supplementary access from NMBU Orion HPC. Access to the LUMI supercomputer (project code NN11074K) was granted by Sigma2; the LUMI supercomputer is owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through Sigma2—Norway. We also thank the editors and reviewers for their insightful comments and constructive feedback.

### v2

@@ 5.1. Suggestions for Future Work
The framework could be extended to video by incorporating temporal frequency analysis. There’s also significant potential in creating adaptive masking strategies that automatically adjust to different generator types and select transforms (e.g., FFT vs. DCT) based on data characteristics. These extensions would build upon the current findings while addressing the identified limitations, moving toward more comprehensive deepfake detection solutions.
@@ Acknowledgements.
This research is supported by the National Research Foundation, Singapore, under its AI Singapore Programmes (AISG Award No.: AISG2-TC-2022-007) and the SUTD project PIE-SGP-AI-2018-01, as well as by the Agency for Science, Technology and Research (A*STAR) under its MTC Programmatic Funds (Grant No. M23L7b0021). Additional support was provided by Changi General Hospital and the Singapore University of Technology and Design under the HealthTech Innovation Fund (HTIF Award No. CGH-SUTD-2021-004). The computational work was supported by resources from Sigma2—the National Infrastructure for High-Performance Computing and Data Storage in Norway—including access to the LUMI supercomputer (project code NN11074K). The LUMI supercomputer is owned by the EuroHPC Joint Undertaking and is hosted by CSC (Finland) and the LUMI consortium through Sigma2—Norway. Supplementary computational support was provided by the Orion High Performance Computing Center (OHPCC) at the Norwegian University of Life Sciences (NMBU). The first author is supported by a PhD scholarship from the Faculty of Science and Technology (REALTEK), Norwegian University of Life Sciences (NMBU). We also thank the editors and reviewers for their insightful comments and constructive feedback.